A standardized exams scores are normally distributed In a recent year, the mean test score was 1495 and the standard deviation was 315. The test scores of four students selected at random are 1900, 1240, 2230, and 1400 Find the z-scores that correspond to each value and determine whether any of the values are unusual The z-score for 1900 is (Round to two decimal places as needed) The Z-score for 1240 is (Round to two decimal places as needed.) The z-score for 2230 is (Round to two decimal places as needed.) The z-score for 1400 is [] (Round to two decimal places as needed.) Which values, if any, are unusual? Select the correct choice below and, if necessary, fillin the answer box within your choice 0 A. The unusual value(s) isare | (Use a comma to separate answers as needed.) O B. None of the values are unusual.

media%2F246%2F24608a0a-474a-43e2-902a-a7
media%2Fe2d%2Fe2d39f41-d3a6-4924-b520-44
Standard Normal Table (Page 1) -3.4 0002 0003 0003 0003 0003 0003 0 0003 0003 0003 -3.3 0003 0004 0004 0004 0004 0004 0004 0005 0005 0005s 3.2 0005 0005 00 0006 0006 0006 0006 0006 0007 0007 3.1 007 0007 0008 0008 0008 0008 0009 0009 0009 0010 3.0 0010 0010 0011 0011 001 0012 0012 0013 0013 0013 -2.9 0014 0014 0015 0015 0016 0016 0017 0018 0018 0019 -2.8 0019 0020 0021 0 0022 0023 023 0024 0025 0026 2.7 0026 0027 0028 0029 0030 0031 0032 0033 0034 0035 -2.6 0036 0037 0038 0039 040 0041 0043 0044 0045 0047 2.4 0064 0066 0068 0069 0071 0073 0075 0078 0080 0082 一2.31.0084 .0087 .0089 .0091 .0094 .0096 .0099 .0102 .0104 .0107 -2.2 0110 0113 0116 011901220125 0129 0132 0136 0139 -2.1 0143 0146 0150 0154 01580162 0166 0170 0174 0179 -2.0 0183 0188 01920197 0202 0207 0212 0217 0222 0228 1.9 0233 0239 0244 0250 0256 02620268 0274 0281 0287 1.8 0294 0301 0307 03140322 0329 0336 0344 0351 035 -1.7 0367 0375 0384 0392 04010409 0418 0427 0436 0446 1.6 0455 0465 0475 0485 0495 05050516 0526 0537 0548 -1.5 0559 0571 0820594 0606 0618 0630 0643 0655 0668 - 1.4 0681 0694 0708 0721 07350749 0764 0778 0793 0808 1.3 0823 0838 0853 086908850901 0918 0934 0951 0968 9 2 2
media%2F822%2F8228c8e1-3025-4ca4-935c-23
media%2F270%2F2703cae0-4c1e-4864-a60b-36
media%2F30f%2F30fb6e0f-75d2-43fb-aab3-00
media%2F384%2F384ce9bc-2c0f-469c-83ad-c2
Answer #1

(a)

\mu = 1495

\sigma = 315

(i)

X = 1900

Z = (X - \mu )/\sigma

= (1900 - 1495)/315 = 1.29

So,

Answer is:

1.29

  (ii)

X = 1240

Z = (X - \mu )/\sigma

= (1240 - 1495)/315 = - 0.81

So,

Answer is:

- 0.81

  (iii)

X = 2230

Z = (X - \mu )/\sigma

= (2230 - 1495)/315 =2.33

So,

Answer is:

2.33

  (iv)

X = 1400

Z = (X - \mu )/\sigma

= (1400 - 1495)/315 = - 0.30

So,

Answer is:

- 0.30

  (b)

Correct option:

A. The unusual value is: 2230

EXPLANATION: X = 2230 has Z score = 2.33 > 2, i.e., X is more than 2 standard deviation from mean.

(c)

Cumulative area = 0.0072

Since the cumulative area is less than 0.5, Z lies on LHS of mid value and is negative.

Standard Normal Table gives Z score corresponding to area = 0.0072 as Z = - 2.445

So

Answer is:

- 2.445